OpenStudy (anonymous):

Let f (x ) = sqrt x and g( x )= x^2 - 6 . Find f (g (x )).

OpenStudy (anonymous):

\[f(g(x))=f(x^2-6)\] is the first step

OpenStudy (anonymous):

then since \(f(\diamondsuit)=\sqrt{\diamondsuit}\) you have \[f(x^2-6)=\sqrt{x^2-6}\]

OpenStudy (anonymous):

oh yay! I got it right =) thank you

OpenStudy (anonymous):

yw